# Turnstile area as a measure for chaotic transport in magnetic confinement fusion devices

### v1

@@ VII Conclusions and Discussion
We have adapted and optimized the calculation for the specific application to three-dimensional magnetic fields, where the integration of field lines is a very costly operation. With the analytical toy model we have shown how different perturbations generate chaotic transport and increase the turnstile area. We have also shown that this calculation is capable of resolving the turnstile area in a very large range of perturbation amplitudes, from very nearly integrable fields, to almost completely stochastic fields.
Transport through an island chain requires a trajectory (field line) to pass through two different turnstiles one after another. The inner turnstile governs passage from the plasma region into the resonance zone, and the outer turnstile then determines passage into the outer chaotic region. If, as in most devices, including the W7-X stellarator, the wall is placed intersecting an edge island, the most important transport channel is through the inner turnstile, as in the island region many trajectories quickly intersect a wall. The calculations of turnstile area in the stellarator configurations show that the outer turnstile area is much larger than the inner turnstile area in the observed configurations. This is likely due to the perturbations due to coil ripple, and the outer manifold passing closer to the coils. Though such fields may look chaotic, the transport from the plasma to the wall is hardly affected by chaos originating from the outer turnstile.
@@
In an optimized stellarator, where neoclassical transport is very low, turbulent transport becomes the dominant mechanism that transports heat perpendicular to the field lines. This can be caused by a wild array of mechanisms and instabilities including drift waves and charge density oscillations, and a predictive theory of turbulent transport in stellarators is nowhere near. Especially in the edge of the stellarator, where density gradients are large, many instabilities exist, that can compete with direct transport along field lines. Transport caused by direct flow of field lines through the turnstile mechanism is therefore not always the most important factor determining heat transport in the stellarator edge. Especially if the turnstile area is small, it is to be expected that neoclassical or turbulent transport can dominate in this region. The calculation of the turnstile area allows for direct quantification of the amount of transport that is due to stochasticity in the magnetic field, and would allow one to potentially to separate these two transport channels in experiment.
@@
The turnstile area also correlates with the stochasticity in the field, and when the turnstile area is zero, the stable and unstable manifold must overlap. The calculation of the turnstile area can be a valuable tool for designing future stellarators, especially when the dynamics on a specific heteroclinic connection (i.e. the inboard or the outboard side of the island) is important for the transport. Using such calculations, the amount of stochasticity in the edge region can be tuned, both on the inner and on the outer turnstile of an island chain. The capability to optimize the turnstile area is in the process of being developed and will be the subject of future publications.
@@ Acknowledgements.
The authors would like to thank Jim Meiss for the inspiring discussions that led to this work about implementing the action calculation in magnetic fields. The authors would like to thank Matt Landreman for invaluable feedback in general, and in particular his suggestions for improving the heteroclinic trajectory finding algorithm. Similarly, our eternal gratitude goes out to Andrew Giuliani for creating and maintaining the QUASR database, which provides a near endless supply of interesting configurations to analyze. This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). The Swiss contribution to this work has been funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.

### v3

@@ VII Conclusions and Discussion
We have adapted and optimized the calculation for the specific application to three-dimensional magnetic fields, where the integration of field lines is a relatively costly operation. With the analytical toy model we have shown how different perturbations generate chaotic transport and increase the turnstile area. We have also shown that this calculation is capable of resolving the turnstile area in a very large range of perturbation amplitudes, from very nearly integrable fields, to almost completely chaotic fields.
Transport through an island chain requires a trajectory (field line) to pass through two different turnstiles one after another. The inner turnstile governs passage from the plasma region into the resonance zone, and the outer turnstile then determines passage into the outer chaotic region. If, as in most devices, including the W7-X stellarator, the PFCs are placed intersecting an edge island, the most important transport channel is likely to be through the inner turnstile, as in the island region many trajectories quickly intersect a PFC. The calculations of turnstile area in the stellarator configurations show that the outer turnstile area is much larger than the inner turnstile area in the observed configurations. This is likely due to the perturbations due to coil ripple, and the outer manifold passing closer to the coils. Though such fields may look chaotic, the transport from the plasma to the PFCs is hardly affected by chaos originating from the outer turnstile.
@@
In an optimized stellarator, where neoclassical transport is very low, turbulent transport becomes the dominant mechanism that transports heat perpendicular to the field lines. Turbulence can be caused by many mechanisms, and a predictive theory of turbulent transport in stellarators is nowhere near. Transport caused by the turnstile mechanism is therefore not always the most important factor determining heat transport in the stellarator edge. Especially if the turnstile area is small, it is to be expected that other mechanisms such as neoclassical or turbulent transport can dominate. The calculation of the turnstile area allows for direct quantification of the amount of transport that is due to chaos in the magnetic field, and would allow one to potentially separate these two transport channels in experiment.
@@
The turnstile area also correlates with the chaos in the field, and when the turnstile area is zero, the stable and unstable manifold must overlap. The calculation of the turnstile area can be a valuable tool for designing future stellarators, especially when the dynamics on a specific heteroclinic connection (i.e. the inboard or the outboard side of the island) is important for the transport. Using such calculations, the amount of chaos in the edge region can be tuned, both on the inner and on the outer turnstile of an island chain. The capability to optimize the turnstile area is in the process of being developed and will be the subject of future publications.
@@ Acknowledgements.
The authors would like to thank Jim Meiss for the inspiring discussions that led to this work about implementing the action calculation in magnetic fields. The authors would like to thank Matt Landreman for invaluable feedback in general, and in particular his suggestions for improving the heteroclinic trajectory finding algorithm. Similarly, our eternal gratitude goes out to Andrew Giuliani for creating and maintaining the QUASR database, which provides a near endless supply of interesting configurations to analyze. R. Davies would like to acknowledge useful guidance from and discussions with Oliver Schmitz, Kelly Garcia and Dieter Boeyaert (University of Wisconsin-Madison) and Sophia Henneberg, Sergei Makarov and Amit Kharwandikar (Max Planck Institute for Plasma Physics, Greifswald. This work was supported by a grant from the Simons Foundation (1013657, JL). This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). The Swiss contribution to this work has been funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.